OpenStudy (anonymous):

A 7.36 g sample of copper is contaminated with a additional 0.51 g sample of zinc. suppose an atomic mass measurement was performed on this sample. What would be the measured atomic mass?

OpenStudy (chmvijay):

it will show 67.94 for 7.36 g of cu has 0.51 gram of zinc as impurities for 63.54(atomic mass) of cu has =0.51*63.54 /7.36 =4.40 g then the atomic mass of Cu will be =63.54+4.40 = 67.94 right